1. 100:5 ::20:X. Determine the value of X.

2. Complete the following ratio: 10:100::1:X.

X=______
Answer #1

loo !5 !! 20. n loo n = 2005 100 n = 100 n = 100/100 x = 1 101 100 :: 1! 10 n = 100x1 son = 100 x = 100/10 x = 10